OpenStudy (anonymous):

Add or subtract 3/x-3 - 5/x-2

OpenStudy (anonymous):

\[\frac{3}{x-3} - \frac{5}{x-2} \implies \frac{3(x-2) - 5(x-3)}{(x-3)(x-2)}\] Solve this..

OpenStudy (anonymous):

I don't understand how to solve it though.